OpenStudy (anonymous):

I need help again. The Walzer's average phone bill increased from $16.65 to $19.98. What is percent of increase?

OpenStudy (dumbcow):

16.65 represents 100% of 16.65 look at the amount of increase 19.98 - 16.65 = 3.33 to make it a percent divide by starting amount and X100 3.33/16.65 = 0.2 *100 = 20%

OpenStudy (anonymous):

increased= 19.98-16.65=3.33 ( 3.33/16.65)=(x/100) cross and mutiplfly

OpenStudy (anonymous):

thanks:)